OpenStudy (anonymous):

If cotangent = 3/4, find the secant.

OpenStudy (anonymous):

cot A = 3/4 = B/P calculate hypotenous \[H = \sqrt{P ^{2} + B ^{2}}\] sec A = H/B

OpenStudy (anonymous):

Using this, I got 5/3, but none of the answer choices offer that. The answer choices are 1/2, 3/5, 2, and 8/5.

OpenStudy (anonymous):

i think there is some problem in question or answer

OpenStudy (anonymous):

I hate when this happens, it happened yesterday too. :P Thanks for the help though, I think I understand better than what the textbook said.